OpenStudy (anonymous):

if a number 774958A96B is to be divisible by 8 and 9. the values of A and B will be?

OpenStudy (kc_kennylau):

find B first, any numbers that is divisible by 8 have the last three digits divisible by 8

OpenStudy (kc_kennylau):

that means 96B is divisible by 8

OpenStudy (kc_kennylau):

B is 0 or 8

OpenStudy (kc_kennylau):

7+7+4+9+5+8+A+9+6+B =55+A+B =55+A or 63+A When B=0, A=8 When B=8, A=0 or A=9